OpenStudy (anonymous):

e^2x-2e^x=3 how do I solve this?

terenzreignz (terenzreignz):

I'd suggest reducing it to something more familiar, perhaps let \[\large u = e^x\] Thus making it \[\large u^2 - 2u = 3\]

OpenStudy (agent0smith):

Let u=e^x

OpenStudy (agent0smith):

Dammit @terenzreignz

terenzreignz (terenzreignz):

I've heard that many a time, @agent0smith :) (in real life too XD )

OpenStudy (anonymous):

ok then what

OpenStudy (agent0smith):

Then solve it the same way you'd solve a quadratic formula, find u. Then go back and use u=e^x to find x.

terenzreignz (terenzreignz):

Then, you solve for the value of u...\[\large u^2 -2u -3=0\] Using any method that you feel comfy with :)

terenzreignz (terenzreignz):

dammit @agent0smith :D

OpenStudy (agent0smith):

haha :D

OpenStudy (anonymous):

all this is to get you to factor this beast with a little practice you can factor \[e^{2x}-e^x-3=0\] with just a mental substitution

OpenStudy (anonymous):

well not really, but you can for \[e^{2x}-2e^x-3=0\]

OpenStudy (anonymous):

(2e^x-1)(e^x+2)

OpenStudy (anonymous):

I mean 3

OpenStudy (anonymous):

careful

OpenStudy (anonymous):

?

terenzreignz (terenzreignz):

Try factoring this first... \[\large u^2 - 2u - 3\]

OpenStudy (anonymous):

oh you pull the 2 out first? so 2(u-1)(u+3)

terenzreignz (terenzreignz):

There is no "2" to pull out...

OpenStudy (anonymous):

i think im confusing myself

terenzreignz (terenzreignz):

Don't give up, though. I'm not really good at explaining how to factor, so you're going to have to get that from someone else...

OpenStudy (anonymous):

I think I got it

terenzreignz (terenzreignz):

What did you get?

OpenStudy (anonymous):

e^x+1=0 so e^x=-1 and then e^x=3 for the other one

terenzreignz (terenzreignz):

Well, can \[\Large e^x=-1\] have a solution? To which exponent do you raise e so that it becomes negative? That's not entirely infeasible, but it's probably beyond the scope of this question.... btw for future reference... \[\huge e^{i\pi}= -1\]

terenzreignz (terenzreignz):

So the only thing you have to worry about is \[\huge e^x = 3\] so what's x?

OpenStudy (anonymous):

so you just take the ln of 3 right?1.0986

terenzreignz (terenzreignz):

Yes. Precisely :)